OpenStudy (anonymous):

If X = 10 cm and Y = 24 cm, what is the length of Z? A. 23 cm B. 25 cm C. 29 cm D. 26 cm

OpenStudy (anonymous):

in what?

OpenStudy (anonymous):

triangle

OpenStudy (anonymous):

What kind of triangle? Is there a picture? All those choices are valid for the 3rd side of a triangle.

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

Right triangle. So, using pythagorean theorem, you have: z^2 = x^2 + y^2 => z^2 = 100 + 576 => z^2 = 676 => z = SQRT(676) => z = 26 cm

OpenStudy (anonymous):

thanks